OpenStudy (anonymous):

while racing on a flat track, a car rounds a curve of 28m radius and instantaneously experiences a centripetal acceleration of 18 m/s2. how fast was the car going? how would you go about doing this?

OpenStudy (aaronq):

you need to use the equations of uniform circular motion (see here: http://www.sparknotes.com/physics/dynamics/uniformcircularmotion/section1.rhtml) for this problem you need \(\sf \large a=\dfrac{v^2}{r}\rightarrow v=\sqrt{ar}\)

OpenStudy (anonymous):

so would that mean that its the square root of 504?

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

thank you :)

OpenStudy (aaronq):

no problem !